OpenStudy (anonymous):

solve for x 4x(x-7)-5x(x-6)=-3 i got x=sqrt5 ??

OpenStudy (saifoo.khan):

Nope, incorrect.

OpenStudy (mimi_x3):

x=-1 or x= 3

OpenStudy (anonymous):

can you show me the steps for getting that?

OpenStudy (mimi_x3):

Saifi can do it :P

OpenStudy (saifoo.khan):

4x^2 - 28x - 5x^2 +30x = -3 -x^2 +2x +3 = 0

OpenStudy (saifoo.khan):

-x^2 +2x +3 = 0 x^2 -2x -3 =0 (x+1)(x-3)

OpenStudy (mimi_x3):

C.Saifi has the stamina to do it from last night :P

OpenStudy (anonymous):

ok, i see where i missed a step lol, thanks!

OpenStudy (saifoo.khan):

np.

OpenStudy (mimi_x3):

Oh gawd, you want people to know what's going on. God bless you Saifi, I feel sorry for you and your gf :P

OpenStudy (saifoo.khan):

Haha, you are the one who started here.

OpenStudy (mimi_x3):

I did NOT, you starteed it first so i continued it, its common sense

OpenStudy (saifoo.khan):

LOL, CS.. you dont have it i know.

OpenStudy (mimi_x3):

How would you know ? You're not me ? I know that you don't have it since your saying it